### 23 Feb 2024 v1

@@ Abstract
Recent years have witnessed great progress in image restoration thanks to the advancements in modern deep neural networks e.g. Convolutional Neural Network and Transformer. However, existing restoration backbones are usually limited due to the inherent local reductive bias or quadratic computational complexity. Recently, Selective Structured State Space Model e.g., Mamba, has shown great potential for long-range dependencies modeling with linear complexity, but it is still under-explored in low-level computer vision. In this work, we introduce a simple but strong benchmark model, named MambaIR, for image restoration. In detail, we propose the Residual State Space Block as the core component, which employs convolution and channel attention to enhance the capabilities of the vanilla Mamba. In this way, our MambaIR takes advantage of local patch recurrence prior as well as channel interaction to produce restoration-specific feature representation. Extensive experiments demonstrate the superiority of our method, for example, MambaIR outperforms Transformer-based baseline SwinIR by up to 0.36dB, using similar computational cost but with a global receptive field. Code is available at https://github.com/csguoh/MambaIR.
@@ 1 Introduction
Image restoration, aiming to reconstruct a high-quality image from a given low-quality input, is a long-standing problem in computer vision and further has a wide range of sub-problems such as super-resolution, image denoising, etc. Since several pioneering works [64, 65, 13], Convolutional Neural Networks (CNNs) have gained great success than traditional model-based method [21, 10], and almost dominate this field in the past few years [30, 56, 35, 67, 11, 63].
@@
Despite the promising results, these restoration backbones usually suffer from the inherent problem of CNNs. First, the weight-sharing strategy in CNN makes the model remain static for varying inputs, preventing robustness on unseen real-world low-quality images. Second, the receptive field of one convolution layer is constrained by the kernel size, which leads to strong local reductive bias. Although stacking multiple convolutional layers can theoretically enlarge the receptive field, however as shown in Fig. 1(a) the effective receptive field [38] of CNN-based restoration methods is still limited.
Inspired by the success in natural language processing, Transformer [54] has proven superior performance than CNNs in several vision tasks due to weak inductive bias as well as global interaction modeling [4, 14, 37]. However, taming Transformer for image restoration poses a significant challenge since the self-attention mechanism scales quadratically with the number of tokens. Considering that image restoration methods usually set the number of tokens to the image resolution, therefore directly using the attention mechanism will incur a huge computational overhead. Although some works make efforts to divide input into small patches [6] or employ shifted window attention [34, 7], these strategies harm the long-term dependency [7] as shown in Fig. 1(b), and do not essentially address the quadratic complexity problem.
Recently, structured state-space sequence models (S4), have emerged as an efficient and effective backbone for constructing deep networks [18, 52, 15, 43, 17]. In particular, the improved S4 with selective mechanism and efficient hardware design, also called Mamba, has been proven to surpass Transformer on tasks requiring long-term dependency modeling such as natural language processing [17]. More recently, some variants of Mamba have also presented promising results in computer vision [72, 36, 40], such as image classification, biomedical image segmentation, etc. However, the power of Mamba in low-level vision is still under-explored, which motivates us to explore the potential of Mamba to enhance the long-range modeling ability for image restoration networks.
In this work, we propose a simple but effective benchmark restoration model, namely MambaIR, to adapt Mamba for image restoration. Following  [34, 7, 62], our MambaIR consists of three stages: shallow feature extraction, deep feature extraction, and high-quality image reconstruction. The shallow feature extraction employs a simple convolution layer to extract the shallow feature. Then the deep feature extraction is performed with several stacked Residual State Space Blocks (RSSBs). As the core component of MambaIR, the RSSB is designed to enhance the capabilities of the original Mamba with image restoration priors. Specifically, to tackle the problem that the original Mamba treats each token equally, ignoring the patch recurrence occurring in the neighborhood, we add additional convolution layers to give local pixels more focus. We further introduce channel attention layer [22] to supplement channel interaction which is beneficial for image restoration but missed in Mamba. In addition, we use the skip connection within each block to offer shortcuts for feature aggregation. Finally, the high-quality image reconstruction stage fuses both shallow and deep features to restore the high-quality output image.
Our main contributions can be summarized as follows:
@@
  • We are the first to work to adapt state space models for low-level image restoration via extensive experiments to formulate MambaIR, which acts as a simple but effective alternative for CNNs and Transformers.
@@
  • We propose the Residue State Space Block (RSSB) which can boost the power of the original Mamba with local spatial prior and channel interaction.
@@
  • Extensive experiments on various tasks demonstrate our MambaIR outperform strong Transformer-based baselines to provide a strong and promising backbone model solution for image restoration.
@@ 2.1 Image Restoration
Image restoration has been significantly advanced since the introduction of deep learning by several pioneering works, such as SRCNN [13] for image super-resolution, DnCNN [64] for image denoising and ARCNN [12] for JPEG compression artifact reduction, etc. Early attempts usually elaborate CNNs with techniques such as residual connection [27, 63, 5], dense connection [56, 69] and others [29, 58, 16, 68, 11] to improve model representation ability. Despite the success, CNN-based restoration methods typically face challenges in effectively modeling global dependencies as well as static weight for varying inputs. As Transformers have been proven to be a strong competitor for CNNs in multiple computer vision tasks [4, 14, 37], using Transformer for image restoration appears promising. However, it still faces specific challenges from the quadratic computational complexity of the self-attention mechanism [54]. To address this, IPT [6] divides one image into several small patches and processes each patch independently with self-attention. SwinIR [34] further introduces shifted window attention to improve the performance. In addition, promising progress continues to be made in designing efficient and effective attention mechanisms for restoration [62, 7, 31, 9, 59, 8]. Nonetheless, efficient attention design usually comes at the expense of global effective receptive fields, and the issue of quadratic computational complexity is not essentially addressed.
@@ 2.2 State Space Models
State Space Models (SSMs) [19, 18, 52], stemming from classics control theory [26], are recently introduced to deep learning as a competitive backbone for state space transforming. The promising property of linearly scaling with sequence length in long-range dependency modeling has attracted great interest from searchers. For example, the Structured State-Space Sequence model (S4) [18] is a pioneer work for the deep state-space model in modeling the long-range dependency. Later, S5 layer [52] is proposed based on S4 and introduces MIMO SSM and efficient parallel scan. Moreover, H3 [15] achieves promising results that nearly fill the performance gap between SSMs and Transformers in natural language.  [43] further improve S4 with gating units to obtain the Gated State Space layer to boost the capability. More recently, Mamba [17], a data-dependent SSM with selective mechanism and efficient hardware design, outperforms Transformers on natural language and enjoys linear scaling with input length. And there are also pioneering works that adopt Mamba to vision tasks such as image classification [36, 72], video understanding [55], biomedical image segmentation [40] and others [25, 47]. In this work, we explore the potential of Mamba to image restoration with restoration-specific prior to sever as a simple but effective baseline for future work.
@@ 3 Methodology
This work aims to introduce the recent advanced state space model (SSM), i.e., Mamba, to low-level image restoration. In this section, we begin with a description of the preliminaries of the SSM. Then we give an overview of our simple but effective benchmark model MambaIR, followed by a detailed description of the proposed Residule State-Space Block (RSSB). Finally, we provide a comprehensive discussion of the difference between our MambaIR and previous methods.
@@ 3.1 Preliminaries
After the discretization, the discretized version of Eq. 1 with step size $\rm\Delta$ can be rewritten as:
@@
However, the formulation in Eq. 3 focuses on the LTI system with parameters remaining static for varying inputs. To address this limitation, recent effort [17] attempt to incorporate the selective scan mechanism in which the matrices $\overline{\rm\textbf{B}}$, C and $\rm\Delta$ are derived from the input data. Inspired by the favorable performance, we also employ the selective scan mechanism to facilitate the awareness of the contextual information embedded in the input.
@@
3.2 Overall Architecture
As shown in Fig. 2, our MambaIR consists of three stages: shallow feature extraction, deep feature extraction, and high-quality reconstruction. Given a low-quality (LQ) input image $I_{LQ}\in\mathbb{R}^{H\times W\times 3}$, we first employ a $3\times 3$ convolution layer from the shallow feature extraction to generate the shallow feature $F_{S}\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ represent the height and width of the input image, and $C$ is the number of channels. Subsequently, the shallow feature $F_{S}$ undergoes the deep feature extraction stage to acquire the deep feature $F_{D}^{l}\in\mathbb{R}^{H\times W\times C}$ at the $l$-th layer, $l\in\{1,2,\cdots L\}$. This stage is stacked by multiple Residual State-Space Groups (RSSGs), with each RSSG containing several Residue State-Space Blocks (RSSBs). Moreover, an additional convolution layer is introduced at the end of each group to refine features extracted from RSSB. Finally, we use the element-wise sum to obtain the input of the high-quality reconstruction stage $F_{R}=F_{L}+F_{S}$, which is used to reconstruct the high-quality (HQ) output image $I_{HQ}$.
@@ 3.3 Residual State-Space Block
The block design in previous Transformer-based image restoration networks [34, 62, 7, 9] mainly follow the Norm $\rightarrow$ Attention $\rightarrow$ Norm $\rightarrow$ MLP flow. Although self-attention and SSM are similar in modeling long-range dependencies, however, we experimentally find that simply replacing Attention with SSM can only obtain sub-optimal results. Therefore, it is promising to tailor a brand-new block structure for Mamba-based restoration networks.
To this end, we propose the Residual State-Space Block (RSSB) to improve the original SSM block with specific restoration priors. Specifically, we consider the local patch recurrence [33] as well as the channel-wise interaction [22] to assist the long-range spatial molding in Mamba. As shown in Fig. 2(a), given the input deep feature $F_{D}^{l}\in\mathbb{R}^{H\times W\times C}$, we first use the Layernorm (LN) followed by the Vision State-Space Module (VSSM) [36] to capture the spatial long-term dependency. Following [8], we also use learnable scale factor $s\in\mathbb{R}^{C}$ in skip connection [8]for better performance:
@@
After that, we employ another LayerNorm to normalize the $Z^{l}$ and then use convolution layers to model the spatial local similarity prior. In order to mitigate channel redundancy as well as efficiency, the convolution layer adapts the bottleneck structure, i.e., the channel is first compressed by a factor $\gamma$ to obtain features with the shape $\mathbb{R}^{H\times W\times\frac{C}{\gamma}}$, then we perform channel expansion to recover the original shape. Furthermore, to model the interaction among channels, we introduce the Channel-Attention layer (CA) [22] followed by another tunable scale factor $s^{\prime}\in\mathbb{R}^{C}$ in residual connection to acquire the final output $F_{D}^{l+1}$ of the RSSB. The above process can be formulated as:
@@ 3.4 Vision State-Space Module
To maintain efficiency, the Transformer-based restoration networks usually divide input into small patches [6] or adopt shifted window attention [34], hindering the interaction at the whole-image level. Motivated by the success of Mamba in long-range modeling with linear complexity, we introduce the Vision State-Space Module to image restoration. Following [36], the input feature $X\in\mathbb{R}^{H\times W\times C}$ will go through two parallel branches. In the first branch, the feature channel is expanded to $\lambda C$ by a linear layer, where $\lambda$ is a pre-defined channel expansion factor, followed by a depth-wise convolution, SiLU [51] activation function, together with the 2D-SSM layer and Layernorm. In the second branch, the features channel is also expanded to $\lambda C$ with a linear layer followed by the SiLU activation function. After that, features from the two branches are aggregated with the Hadamard product. Finally, the channel number is projected back to $C$ to generate output $X_{out}$ with the same shape as input.
@@
The standard Mamba [17] causally processes the input data, and thus can only capture information within the scanned part of the data. This property is well suited for NLP tasks that involve a sequential nature but poses significant challenges when transferring to non-causal data such as images. To better utilize the 2D spatial information, we follow [36] and introduce the 2D Selective Scan Module (2D-SSM). As shown in Fig. 2(c), the 2D image feature is flattened into a 1D sequence with scanning along four different directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right. Then the long-range dependency of each sequence is captured according to the discrete state-space equation in Eq. 3. Finally, all sequences are merged using summation followed by the reshape operation to recover the 2D structure.
3.6 Discussion
Here we would like to discuss the differences between our MambaIR with previous image restoration arts to highlight the contribution of this work. Our MambaIR improves the recent advanced state-space model Mamba with restoration-specific priors, thus allowing the injection of local patch recurrence and channel interaction while maintaining strong long-range capability. Different from Transformer-based methods, the sequential causal property of SSMs exempts us from positional encoding [54]. Meanwhile, the linear complexity with input scale ensures the training and inference efficiency, thus avoiding the shifted window operation which may induce an unsatisfactory effective receptive field.

### 25 Mar 2024 v2

@@ Abstract
Recent years have seen significant advancements in image restoration, largely attributed to the development of modern deep neural networks, such as CNNs and Transformers. However, existing restoration backbones often face the dilemma between global receptive fields and efficient computation, hindering their application in practice. Recently, the Selective Structured State Space Model, especially the improved version Mamba, has shown great potential for long-range dependency modeling with linear complexity, which offers a way to resolve the above dilemma. However, the standard Mamba still faces certain challenges in low-level vision such as local pixel forgetting and channel redundancy. In this work, we introduce a simple but effective baseline, named MambaIR, which introduces both local enhancement and channel attention to improve the vanilla Mamba. In this way, our MambaIR takes advantage of the local pixel similarity and reduces the channel redundancy. Extensive experiments demonstrate the superiority of our method, for example, MambaIR outperforms SwinIR by up to 0.45dB on image SR, using similar computational cost but with a global receptive field. Code is available at https://github.com/csguoh/MambaIR.
@@ 1 Introduction
Image restoration, aiming to reconstruct a high-quality image from a given low-quality input, is a long-standing problem in computer vision and further has a wide range of sub-problems such as super-resolution, image denoising, etc. With the introduction of modern deep learning models such as CNNs [70, 14, 37, 76, 12] and Transformers [8, 36, 9, 34, 11], state-of-the-art performance has continued to be refreshed in the past few years.
To some extent, the increasing performance of deep restoration models largely stems from the increasing network receptive field. First, a large receptive field allows the network to capture information from a wider region, enabling it to refer to more pixels to facilitate the reconstruction of the anchor pixel. Second, with a larger receptive field, the restoration network can extract higher-level patterns and structures in the image, which can be crucial for some structure preservation tasks such as image denoising. Finally, Transformer-based restoration methods which possess larger receptive fields experimentally outperform CNN-based methods, and the recent art [9] also points out that activating more pixels usually leads to better restoration results.
@@
Despite possessing many attractive properties, it appears that there exists an inherent choice dilemma between global receptive fields and efficient computation for current image restoration backbones. For CNN-based restoration networks [37, 76], although the effective receptive field is limited (as shown in Fig. 1(a)), it is appropriate for resource-constrained device deployments due to the favorable efficiency of convolution parallel operations. By contrast, Transformer-based image restoration methods usually set the number of tokens to the image resolution [36, 8, 9], therefore, despite the global receptive field, directly using the standard Transformer [57] will come at an unacceptable quadratic computational complexity. Moreover, employing some efficient attention techniques such as shifted window attention [39] for image restoration, usually comes at the expense of a globally effective receptive field (as shown in  Fig. 1(b)), and does not intrinsically escape out of the trade-off between a global receptive field and efficient computation.
Recently, structured state-space sequence models (S4), especially the improved version Mamba, have emerged as an efficient and effective backbone for constructing deep networks [21, 55, 17, 46, 19]. This development hints at a potential resolution to balancing global receptive field and computational efficiency in image restoration. In detail, the discretized state space equations in Mamba can be formalized into a recursive form and can model very long-range dependencies when equipped with specially designed structured reparameterization [20]. This means that Mamba-based restoration networks can naturally activate more pixels, thus improving the reconstruction quality. Furthermore, the parallel scan algorithm [19] renders Mamba to process each token in a parallel fashion, facilitating efficient training on modern hardware such as GPU. The above promising properties motivate us to explore the potential of Mamba to achieve efficient long-range modeling for image restoration networks.
However, the standard Mamba [19], which is designed for 1D sequential data in NLP, is not a natural fit for image restoration scenarios. First, since Mamba processes flattened 1D image sequences in a recursive manner, it can result in spatially close pixels being found at very distant locations in the flattened sequences, resulting in the problem of local pixel forgetting. Second, due to the requirement to memorize the long sequence dependencies, the number of hidden states in the state space equations is typically large, which can lead to channel redundancy, thus hindering the learning of critical channel representations.
To address the above challenges, we introduce MambaIR, a simple but very effective benchmark model, to adapt Mamba for image restoration. MambaIR is formulated with three principal stages. Specifically, the 1)Shallow Feature Extraction stage employs a simple convolution layer to extract the shallow feature. Then the 2)Deep Feature Extraction stage performs with several stacked Residual State Space Blocks (RSSBs). As the core component of our MambaIR, the RSSB is designed with local convolution to mitigate local pixel forgetting when applying the vanilla Mamba to 2D images, and it is also equipped with channel attention to reduce channel redundancy caused by the excessive hidden state number. We also employ the learnable factor to control the skip connection within each RSSB. Finally, the 3)High-Quality Image Reconstruction stage aggregates both shallow and deep features to produce a high-quality output image. Through possessing both a global effective receptive field as well as linear computational complexity, our MambaIR serves as a new alternative for image restoration backbones.
In short, our main contributions can be summarized as follows:
@@
  • We are the first to work to adapt state space models for low-level image restoration via extensive experiments to formulate MambaIR, which acts as a simple but effective alternative for CNN- and Transformer-based methods.
@@
  • We propose the Residue State Space Block (RSSB) which can boost the power of the standard Mamba with local enhancement and channel redundancy reduction.
@@
  • Extensive experiments on various tasks demonstrate our MambaIR outperforms other strong baselines to provide a powerful and promising backbone solution for image restoration.
@@ 2.1 Image Restoration
Image restoration has been significantly advanced since the introduction of deep learning by several pioneering works, such as SRCNN [14] for image super-resolution, DnCNN [70] for image denoising, ARCNN [13] for JPEG compression artifact reduction, etc. Early attempts usually elaborate CNNs with techniques such as residual connection [29, 67, 6], dense connection [60, 78] and others [31, 62, 18, 76, 12] to improve model representation ability. Despite the success, CNN-based restoration methods typically face challenges in effectively modeling global dependencies. As Transformers have been proven to be a strong competitor for CNNs in multiple computer vision tasks [5, 15, 39], using Transformers for image restoration appears promising. Despite the global receptive field, it still faces specific challenges from the quadratic computational complexity of the self-attention mechanism [57]. To address this, IPT [8] divides one image into several small patches and processes each patch independently with self-attention. SwinIR [36] further introduces shifted window attention [39] to improve the performance. In addition, progress continues to be made in designing efficient attention mechanisms for restoration [66, 9, 32, 11, 63, 10]. Nonetheless, efficient attention design usually comes at the expense of global effective receptive fields, and the dilemma of the trade-off between efficient computation and global modeling is not essentially resolved.
@@ 2.2 State Space Models
State Space Models (SSMs) [22, 21, 55], stemming from classics control theory [28], are recently introduced to deep learning as a competitive backbone for state space transforming. The promising property of linearly scaling with sequence length in long-range dependency modeling has attracted great interest from searchers. For example, the Structured State-Space Sequence model (S4) [21] is a pioneer work for the deep state-space model in modeling the long-range dependency. Later, S5 layer [55] is proposed based on S4 and introduces MIMO SSM and efficient parallel scan. Moreover, H3 [17] achieves promising results that nearly fill the performance gap between SSMs and Transformers in natural language.  [46] further improve S4 with gating units to obtain the Gated State Space layer to boost the capability. More recently, Mamba [19], a data-dependent SSM with selective mechanism and efficient hardware design, outperforms Transformers on natural language and enjoys linear scaling with input length. Moreover, there are also pioneering works that adopt Mamba to vision tasks such as image classification [38, 81], video understanding [58], biomedical image segmentation [42] and others [26, 50]. In this work, we explore the potential of Mamba to image restoration with restoration-specific designs to serve as a simple but effective baseline for future work.
@@ 3.1 Preliminaries
After the discretization, the discretized version of Eq. 1 with step size $\rm\Delta$ can be rewritten in the following RNN form:
@@
Furthermore, the Eq. 3 can also be mathematically equivalently transformed into the following CNN form:
$\overline{\rm\textbf{K}}$ $\triangleq(\mathrm{\textbf{C}}\overline{\mathrm{\textbf{B}}},{\rm{\textbf{C}}}\overline{\rm\textbf{A}}\overline{\rm\textbf{B}},\cdots,{\rm{\textbf{C}}}{\overline{\rm\textbf{A}}}^{L-1}\overline{\rm\textbf{B}}),$ (4)
y $={\rm\textbf{x}}\circledast\overline{\rm\textbf{K}},$
[…]
The recent advanced state-space model, Mamba [19], have further improved $\overline{\rm\textbf{B}}$, C and $\rm\Delta$ to be input-dependent, thus allowing for a dynamic feature representation. The intuition of Mamba for image restoration lies in its development on the advantages of S4 model. Specifically, Mamba shares the same recursive form of  Eq. 3, which enables the model to memorize ultra-long sequences so that more pixels can be activated to aid restoration. At the same time, the parallel scan algorithm [19] allows Mamba to enjoy the same advantages of parallel processing as Eq. 4, thus facilitating efficient training.
3.2 Overall Architecture
As shown in Fig. 2, our MambaIR consists of three stages: shallow feature extraction, deep feature extraction, and high-quality reconstruction. Given a low-quality (LQ) input image $I_{LQ}\in\mathbb{R}^{H\times W\times 3}$, we first employ a $3\times 3$ convolution layer from the shallow feature extraction to generate the shallow feature $F_{S}\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ represent the height and width of the input image, and $C$ is the number of channels. Subsequently, the shallow feature $F_{S}$ undergoes the deep feature extraction stage to acquire the deep feature $F_{D}^{l}\in\mathbb{R}^{H\times W\times C}$ at the $l$-th layer, $l\in\{1,2,\cdots L\}$. This stage is stacked by multiple Residual State-Space Groups (RSSGs), with each RSSG containing several Residue State-Space Blocks (RSSBs). Moreover, an additional convolution layer is introduced at the end of each group to refine features extracted from RSSB. Finally, we use the element-wise sum to obtain the input of the high-quality reconstruction stage $F_{R}=F^{L}_{D}+F_{S}$, which is used to reconstruct the high-quality (HQ) output image $I_{HQ}$.
@@ 3.3 Residual State-Space Block
The block design in previous Transformer-based restoration networks [36, 66, 9, 11] mainly follow the Norm $\rightarrow$ Attention $\rightarrow$ Norm $\rightarrow$ MLP flow. Although Attention and SSM can both model global dependencies, however, we find these two modules behave differently (see supplementary material for more details) and simply replacing Attention with SSM can only obtain sub-optimal results. Therefore, it is promising to tailor a brand-new block structure for Mamba-based restoration networks.
To this end, we propose the Residual State-Space Block (RSSB) to adapt the SSM block for restoration. As shown in Fig. 2(a), given the input deep feature $F_{D}^{l}\in\mathbb{R}^{H\times W\times C}$, we first use the LayerNorm (LN) followed by the Vision State-Space Module (VSSM) [38] to capture the spatial long-term dependency. Moreover, we also use learnable scale factor $s\in\mathbb{R}^{C}$ to control the information from skip connection:
@@
Furthermore, since SSMs process flattened feature maps as 1D token sequences, the number of neighborhood pixels in the sequence is greatly influenced by the flattening strategy. For example, when employing the four-direction unfolding strategy of  [38], only four nearest neighbors are available to the anchor pixel (see  Fig. 3(a)), i.e., some spatially close pixels in 2D feature map are instead distant from each other in the 1D token sequence, and this over-distance can lead to local pixel forgetting. To this end, we introduce an additional local convolution after VSSM to help restore the neighborhood similarity. Specifically, we employ LayerNorm to first normalize the $Z^{l}$ and then use convolution layers to compensate for local features. In order to maintain efficiency, the convolution layer adopts the bottleneck structure, i.e., the channel is first compressed by a factor $\gamma$ to obtain features with the shape $\mathbb{R}^{H\times W\times\frac{C}{\gamma}}$, then we perform channel expansion to recover the original shape.
In addition, SSMs typically introduce a larger number of hidden states to memorize very long-range dependencies, and we visualize the activation results for different channels in  Fig. 3(b) and find notable channel redundancy. To facilitate the expressive power of different channels, we introduce the Channel Attention (CA) [23] to RSSB. In this way, SSMs can focus on learning diverse channel representations after which the critical channels are selected by subsequent channel attention, thus avoiding channel redundancy. At last, another tunable scale factor $s^{\prime}\in\mathbb{R}^{C}$ is used in residual connection to acquire the final output $F_{D}^{l+1}$ of the RSSB. The above process can be formulated as:
@@
Figure 3: (a) Without using local enhancement will cause spatially close pixels (area in the red box) get forgotten in the flattened 1D sequence due to the long distance. (b) We use RELU and global average pooling on the VSSM outputs from the last layer to get the channel activation values. Most channels are not activated (i.e., channel redundancy) when channel attention is not used.
@@ 3.4 Vision State-Space Module
To maintain efficiency, the Transformer-based restoration networks usually divide input into small patches [8] or adopt shifted window attention [36], hindering the interaction at the whole-image level. Motivated by the success of Mamba in long-range modeling with linear complexity, we introduce the Vision State-Space Module to image restoration.
The Vision State-Space Module (VSSM) can capture long-range dependencies with the state space equation, and the architecture of VSSM is shown in  Fig. 2(b). Following [38], the input feature $X\in\mathbb{R}^{H\times W\times C}$ will go through two parallel branches. In the first branch, the feature channel is expanded to $\lambda C$ by a linear layer, where $\lambda$ is a pre-defined channel expansion factor, followed by a depth-wise convolution, SiLU [54] activation function, together with the 2D-SSM layer and LayerNorm. In the second branch, the features channel is also expanded to $\lambda C$ with a linear layer followed by the SiLU activation function. After that, features from the two branches are aggregated with the Hadamard product. Finally, the channel number is projected back to $C$ to generate output $X_{out}$ with the same shape as input:
@@
The standard Mamba [19] causally processes the input data, and thus can only capture information within the scanned part of the data. This property is well suited for NLP tasks that involve a sequential nature but poses significant challenges when transferring to non-causal data such as images. To better utilize the 2D spatial information, we follow [38] and introduce the 2D Selective Scan Module (2D-SSM). As shown in Fig. 2(c), the 2D image feature is flattened into a 1D sequence with scanning along four different directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right. Then the long-range dependency of each sequence is captured according to the discrete state-space equation. Finally, all sequences are merged using summation followed by the reshape operation to recover the 2D structure.
3.6 Loss Function
To make a fair comparison with previous works [76, 36, 66], we optimize our MambaIR with $L_{1}$ loss for image SR, which can be formulated as:
$\mathcal{L}=||I_{HQ}-I_{LQ}||_{1},$ (8)
where $||\cdot||_{1}$ denotes the $L_{1}$ norm. For image denoising, we utilize the Charbonnier loss [7] with $\epsilon=10^{-3}$:
$\mathcal{L}=\sqrt{||I_{HQ}-I_{LQ}||^{2}+\epsilon^{2}}.$ (9)